OpenStudy (anonymous):

Solve the absolute value equation. Check your solutions. | 5x + 13| = –7

OpenStudy (anonymous):

Make two cases: one where \(5x+13 \ge 0\) and one where \(5x + 13 < 0\) and solve each of those cases independently.

OpenStudy (anonymous):

5x+13=-7 5x=-20 x=-4 OR 5x-13=-7 5x=6 x=6/5 Check by substituting in.

OpenStudy (campbell_st):

the absolute value of 5x + 13 is always positive I'd say it can't equal -7

OpenStudy (anonymous):

5x + 13 = -7 5x = -7 -13 5x = -20 divide both sides by 5. x=-4 :)

OpenStudy (campbell_st):

its interesting x = -4 | 5*(-4) + 13| = | -7| = 7 which doesn't equal -7

OpenStudy (campbell_st):

and | 5*6/5 + 13| = | 19| = 19 and not -7

OpenStudy (campbell_st):

Oh hold on Some1 made a mistake with his 2nd solution -5x - 13 = -7 -5x = 6 x = -6/5 lets check it | 5*-6/5 + 13| = | 7| = 7 so that doesn't equal - 7 either...

OpenStudy (anonymous):

awww, okay lets try and get it right.